# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ Abstract.
We derive the quantitative propagation of chaos in the sense of relative entropy for the 2D viscous vortex model with general circulations, approximating the vorticity formulation of the 2D Navier-Stokes equation on the whole Euclidean space, which is an extension of the previous works [FHM14, JW18]. We improve our previous results in [FW23] to the more general case that the vorticity may change sign, allowing the circulations to be in different orientations. We also extend the results of [Wan24] to the whole space to obtain the sharp local propagation of chaos in the high viscosity regime. Both results of convergence rates are global-in-time on any finite time horizon and optimal among existing literature, thanks to the careful estimates using the Grigor’yan parabolic maximum principle and a new ODE hierarchy and iterated integral estimates.
@@ 1. Introduction
Consider the $N$-particle point vortex system positioning on the whole Euclidean space ${\mathbb{R}}^{2}$ with general circulations for vortices, whose evolution is governed by the following stochastic differential equations (SDEs):
@@
The particles are labeled by the index $i$ ranging from 11 11 to $N$, with $X_{i}(t)\in{\mathbb{R}}^{2}$ and ${\mathcal{M}}_{i}\in{\mathbb{R}}$ representing respectively the position and the circulation of the $i$-th vortex. Here $\{B_{i}(t)\}_{1\leq i\leq N}$ are $N$ independent copies of the standard 2D Brownian motion. The constant $\sigma>0$ in the diffusion term represents the viscosity of the dynamics of the system.
As we can see from our SDE system (1.1), the evolution of the position of each particle is driven by a deterministic drift term, in the form of pairwise interacting force with some weights given by the circulations of other particles, and an extra stochastic diffusion term given by the additive noise with constant strength. The different cases of circulations with ${\mathcal{M}}_{i}>0$ and ${\mathcal{M}}_{i}<0$ represent the two different orientations of the point vortices. We mention that the scale $1/N$ appearing in the vortex circulation is in the sense of mean-field scaling, keeping the whole circulation of the system bounded.
The singular interacting kernel $K$ has many different options arising from various natural physical models. In our study, this kernel is given by the famous 2D Biot-Savart law in fluid mechanics,
@@
where $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$. While the kernel has some singularity of $O(1/|x|)$ at the origin, making the interaction between two particles large enough when they become sufficiently close to each other, previous work by Osada [Osa86] has shown that with probability one the particles actually never collide in the evolution process. Hence the singularity of the system will never be reached, which allows us to use the convention that $K(0)=0$, and thus add the terms with $i=j$ in the summation of (1.1). We also refer the well-posedness and first partial propagation of chaos results of the SDEs (1.1) to [Osa86]. The key regularity property we shall invoke in this article is that $K\in W^{-1,\infty}$, with $K=\nabla\cdot V$ as in [JW18], where
$V(x)=\Big{(}-\frac{1}{2\pi}\arctan\frac{x_{1}}{x_{2}}\Big{)}\text{Id}$
is bounded on the whole space, with $\|V\|_{\infty}=1/4$.
Throughout this article, we make the common assumption that ${\mathcal{M}}_{i}$ is independent of time $t$, which means that the circulation of each vortex, once being set up at the initial time $t=0$, keeps unchanged as the system evolves in time. We also assume that $\{{\mathcal{M}}_{i}\}_{1\leq i\leq N}$ are $N$ independent copies of some random variable ${\mathcal{M}}$ compactly supported in $[-A,A]$ for some $A>0$. Typical examples include the uniform distribution $U(-A,A)$.
@@
Now the new point vortex system (1.2) is endowed with scaled circulations ~$\tilde{{\mathcal{M}}}_{j}={\mathcal{M}}_{j}/A$, which are supported in $[-1,1]$. Hence we can always change our target system into the assumption case $A\leq 1$, under the price that the viscosity constant $\sigma$ is rescaled by the constant $A$ as well, which is consistent with our assumptions on $\sigma$ in the main results.
Our assumption made on ${\mathcal{M}}_{i}$ corresponds to the practical statistical procedure that we put some i.i.d. circulations on each point vortex to expect the convergence of the vorticity as $N\rightarrow\infty$ in some Law of Large Numbers sense. This setting is a generalized version of the previous work by the authors [FW23], where the point vortex system has fixed circulations ${\mathcal{M}}_{i}\equiv 1$ at each vortex $X_{i}$ and thus the particle system is symmetric in position $X_{i}$. Note that this is not the case in this article, since now we can only assume symmetry on the full information $Z_{i}=({\mathcal{M}}_{i},X_{i})$ for each particle.
We study the mean-field limit dynamics of the above point vortex system (1.1) when $N\rightarrow\infty$ in some appropriate sense of statistical mechanics. The limit self-consistent McKean-Vlasov system writes
@@
represents the joint density of ${\mathcal{M}}$ and $X_{t}$. By Itô’s formula, it is straightforward to check that the vorticity $\omega_{t}$ satisfies the aggregation-diffusion type equation, which is exactly the well-known vorticity formulation of the 2D Navier-Stokes equation in fluid mechanics,
@@
The relation between the particle system (1.1) and the limit McKean-Vlasov system (1.3) as $N\rightarrow\infty$ lies in the regime of the mean-field theory and is described by the so-called propagation of chaos argument. For our stochastic point vortex system approximating the limit 2D Navier-Stokes equation, propagation of chaos in $L^{1}$ sense can be found firstly in Fournier-Hauray-Mischler [FHM14], where the authors studied the entropy and Fisher information of the point vortex system, using a compactness argument to derive propagation of chaos both in the entropy and the $L^{1}$ sense, but without any explicit rate. For detailed information and comprehensive results of rencet development of mean-field limit and propagation of chaos, we refer to the reviews such as [Szn91, Jab14, Gol16, JW17].
@@
A special case of our stochastic point vortex system (1.1), say the fixed circulations case ${\mathcal{M}}_{i}\equiv 1$, has been well-studied for approximating the 2D Navier-Stokes equation (1.4), see for instance the previous quantitative propagation of chaos results [JW18, GLBM24, FW23]. However, the simplified model is restricted to the case that all circulations have the same sign and magnitude, thus cannot deal with the special vorticity that $\int_{{\mathbb{R}}^{2}}\omega_{t}\,\mathrm{d}x=0$ for example. This is also reflected by the parabolic maximum principle that $\inf\omega_{t}\geq\inf\omega_{0}=\inf g_{0}\geq 0$, which yields that the vorticity cannot change sign. We view it as an essential limitation for physical considerations, where one may generally face the $L^{1}$ vorticity function, while in this article we resolve this difficulty by allowing the circulations to be positive at some points and negative elsewhere.
1.1. The Liouville Equation, BBGKY Hierarchy and the Limit Equation
In this subsection, we present the corresponding Liouville (master) equation (forward Kolmogorov equation) and the BBGKY hierarchy of the SDE system (1.1) and the nonlinear Fokker-Planck equation of the limit system (1.3), since our method will always work on the PDE level. Hereinafter we keep the following notations
${\mathcal{M}}^{N}=(m_{1},...,m_{N})\in{\mathbb{R}}^{N},\quad X^{N}=(x_{1},...,x_{N})\in{\mathbb{R}}^{2N},\quad Z^{N}=({\mathcal{M}}^{N},X^{N}),\quad z=(m,x).$
Denote the joint distribution of ${\mathcal{M}}^{N}$ and $X^{N}$ for the particle system at time $t$ by $G_{t}^{N}({\mathcal{M}}^{N},X^{N})$, with the tensorized initial data
$G_{0}^{N}(Z^{N})=g_{0}^{\otimes N}(Z^{N})=\prod_{i=1}^{N}g_{0}(z_{i}).$
And we also define the $k$-marginal of the joint law of the particle system as
$G_{t}^{N,k}(z_{1},...,z_{k})=\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N-k}}G_{t}^{N}(Z^{N})\,\mathrm{d}z_{k+1}...\,\mathrm{d}z_{N}.$
@@ 1.1. The Liouville Equation, BBGKY Hierarchy and the Limit Equation
We can relax the i.i.d. assumption on the initial data $G_{0}^{N}$. For our purposes below, it is sufficient to assume that the marginals of all orders of the initial distribution $G_{0}^{N,k}$ are close to their tensorized counterparts $g_{0}^{\otimes k}$ in the sense of relative entropy as below:
$H(G_{0}^{N,k}|g_{0}^{\otimes k})=\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G_{0}^{k}\log\frac{G_{0}^{k}}{g_{0}^{\otimes k}}\,\mathrm{d}Z^{k}\lesssim\frac{k^{2}}{N^{2}}.$
We choose the simplest i.i.d. assumption here, which corresponds to that the relative entropy $H(G_{0}^{N,k}|g_{0}^{\otimes k})$ vanishes, but our results can be easily applied to the above more general case, for instance the initial data $G_{0}^{N}$ can be chosen as in a family of the Gibbs measures.
Applying Itô’s formula to the particle system (1.1), one concludes that the joint law $G_{t}^{N}$ solves the following Liouville equation
@@
We emphasize that $G_{t}^{N}$ shall be understood as the entropy solution of the Liouville equation (1.5). The definition of entropy solution is given as in Jabin-Wang [JW18]. For the sake of completeness, we state it as follows.
@@
A probability density function $G_{t}^{N}\in L^{1}(({\mathbb{R}}\times{\mathbb{R}}^{2})^{N})$ with $G_{t}^{N}\geq 0$ and (×$\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G_{t}^{N}\,\mathrm{d}Z^{N}=1$ is called an entropy solution to the Liouville equation (1.5) in $[0,T]$, if $G_{t}^{N}$ solves (1.5) in the sense of distribution and for a.e. $t\in[0,T]$ it holds that
(×$\,\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G_{t}^{N}\log G_{t}^{N}\,\mathrm{d}Z^{N}+\sigma\sum_{i=1}^{N}\int_{0}^{t}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\frac{|\nabla_{x_{i}}G_{s}^{N}|^{2}}{G_{s}^{N}}\,\mathrm{d}Z^{N}\,\mathrm{d}s$
$\leq$ (×$\,\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G_{0}^{N}\log G_{0}^{N}\,\mathrm{d}Z^{N}.$
For the existence of such entropy solution, we also refer to [JW18].
@@
$+\frac{N-k}{N}\sum_{i=1}^{k}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G_{t}^{N,k+1}\,\mathrm{d}z_{k+1}=\sum_{i=1}^{k}\Delta_{x_{i}}G_{t}^{N,k}.$
Hereinafter we assume that $1\leq k\leq N$ and use the convention that $G_{t}^{N,N+1}=0$. The above equation (1.6) describes the evolution of the $k$-marginal using the next order $(k+1)$-marginal, hence forms a hierarchy of PDEs, which is indeed the BBGKY hierarchy of the particle system.
As for the limit system (1.3), by applying Itô’s formula, we also obtain that the limit density $g_{t}(m,x)$ solves the following nonlinear Fokker-Planck equation
@@
After direct and simple computations, one can verify that the $k$-tensorized law of $g_{t}(m,x)$, denoted by $g_{t}^{k}({\mathcal{M}}^{k},X^{k})$, solves the following PDE:
@@
Finally we define our normalized relative entropy between the joint distribution of the particle system and the $k$-tensorized law of the limit system, which will be the major quantity that we shall estimate.
@@
The normalized relative entropy between two probability measures $F_{k}$ and $f_{k}$ on ${\mathbb{R}}^{k}\times{\mathbb{R}}^{2k}$ is defined by
$H_{k}(F_{k}|f_{k})=\frac{1}{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}F_{k}\log\frac{F_{k}}{f_{k}}\,\mathrm{d}Z^{k}.$
Hereinafter we generally drop the index $t$ for simplicity. Also, due to the fact that the variable ${\mathcal{M}}$ is independent of time $t$, if we let $f^{m}(x)=g(m,x)=\mathbb{P}(X|{\mathcal{M}}=m)$ be the conditional probability density of $X$ under ${\mathcal{M}}=m$ as in [FHM14], we obtain that $f^{m}$ satisfies the same evolution equation (1.7) with $g$. Hence we shall not make difference between $g(m,x)$ and $f^{m}(x)$ for calculations, and we assume that all derivatives of $g$ without index only act on the $x$ variable.

### v2

@@ Abstract.
We derive quantitative propagation of chaos in the sense of relative entropy for the 2D viscous vortex model with general circulations, approximating the vorticity formulation of the 2D Navier-Stokes equation on the whole Euclidean space, which is an extension of the previous works [17, 30]. We improve our previous results in [16] to the more general case that the vorticity may change sign, allowing the circulations to be in different orientations. We also extend the results of [55] to the whole space to obtain the sharp local propagation of chaos in the high viscosity regime. Both results of convergence rates are global-in-time on any finite time horizon and optimal among existing literature, thanks to the sharp logarithmic growth estimates and a new ODE hierarchy and iterated integral estimates.
@@ 1. Introduction
Consider the $N$-particle stochastic point vortex system positioning on the whole Euclidean space ${\mathbb{R}}^{2}$ with general circulations for vortices, whose evolution is governed by the following stochastic differential equations (SDEs):
@@
The particles are labeled with the index $i$ ranging from 11 11 to $N$, with $X_{i}(t)\in{\mathbb{R}}^{2}$ and ${\mathcal{M}}_{i}/N\in{\mathbb{R}}$ representing, respectively, the position and circulation of the $i$-th vortex.
The displacement of the particles is determined by the combination effect of two factors. The first one is given by the pairwise interaction force with some weights given by the circulations of other particles. The different cases of circulations with ${\mathcal{M}}_{i}>0$ and ${\mathcal{M}}_{i}<0$ represent the two different orientations of the point vortices. We mention that the scale $1/N$ that appears in the vortex circulation is in the sense of mean-field scaling, keeping the whole circulation of the system bounded. The second one is given by the extra additive noise, driven by $N$ independent copies of the standard 2D Brownian motion, which are denoted by $\{B_{i}(t)\}_{1\leq i\leq N}$. The constant $\sigma>0$ represents the viscosity of the system dynamics and measures the intensity of the diffusion.
The interacting kernel $K$ has many different options arising from various natural physical models. In our study, this kernel is given by the famous 2D Biot-Savart law in fluid mechanics,
@@
where $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$. This point vortex model was initiated in the study of Helmholtz [25] and Kirchhoff [33]. It was Osada [43] who first showed that, although the kernel has some singularity of $O(1/|x|)$ at the origin, which makes the interaction between two particles large enough when they become sufficiently close to each other, the particles actually never collide in the evolution process with probability one. Therefore, the singularity of the system will never be reached, and the well-posedness of (1.1) is justified. This allows us to use the convention that $K(0)=0$, and thus add the terms with $i=j$ to the summation in (1.1).
Throughout this article, we make the common assumption that ${\mathcal{M}}_{i}$ is independent of time $t$, which means that the circulation of each point vortex, once determined at the initial time $t=0$, remains unchanged as the system evolves in time. We also assume that $\{{\mathcal{M}}_{i}\}_{1\leq i\leq N}$ are $N$ independent copies of some random variable ${\mathcal{M}}$ compactly supported in $[-A,A]$ for some $A>0$. Typical examples include the uniform distribution $U(-A,A)$.
@@
This new point vortex system (1.2) is endowed with scaled circulations ~$\tilde{{\mathcal{M}}}_{j}={\mathcal{M}}_{j}/A$, which are supported in $[-1,1]$. Therefore, we can always change our target system to satisfy the additional assumption $A\leq 1$, at the price that the viscosity constant $\sigma$ is also rescaled by the constant $A$, which is consistent with our assumptions on $\sigma$ in the main results.
Our assumption made on ${\mathcal{M}}_{i}$ corresponds to the practical statistical procedure that we put some i.i.d. circulations on each point vortex to expect the convergence of the vorticity as $N\rightarrow\infty$ in some Law of Large Numbers sense. This setting is a generalized version of the previous work of the authors [16], where the point vortex system has fixed circulations ${\mathcal{M}}_{i}\equiv 1$ at each vortex $X_{i}$ and thus the particle system is symmetric in position $X_{i}$. Note that this is not the case in this article, since now we can only assume symmetry on the full information $Z_{i}=({\mathcal{M}}_{i},X_{i})$ for each particle.
We are interested in the mean-field limit dynamics of the point vortex system (1.1) when $N\rightarrow\infty$ in some appropriate sense of statistical mechanics. It is well-known that the limit self-consistent McKean-Vlasov system writes
@@
[…]
@@
The relation between the particle system (1.1) and the limit McKean-Vlasov system (1.3) as $N\rightarrow\infty$ lies in the regime of the mean-field theory and is described by the so-called propagation of chaos argument. For our stochastic point vortex system approximating the limit 2D Navier-Stokes equation, the first such result dates back to Osada [43], under the assumption of large viscosity. Propagation of chaos in the $L^{1}$ sense with positive viscosity was first established in the seminal paper of Fournier-Hauray-Mischler [17], where they studied the entropy and Fisher information of the point vortex system and used a compactness argument, but without any explicit convergence rate. This article is devoted to quantitative propagation of chaos for the point vortex system (1.1). For detailed information and comprehensive results of recent developments in mean-field theory and propagation of chaos, we refer to reviews such as [53, 27, 18, 29].
@@
A special case of our stochastic point vortex system (1.1), say the fixed circulations case ${\mathcal{M}}_{i}\equiv 1$, has been well-studied for approximating the 2D Navier-Stokes equation (1.4), see for instance the previous quantitative propagation of chaos results [30, 22, 16, 55]. However, the simplified model is restricted to the case that all circulations have the same sign and magnitude, thus cannot deal with the special vorticity that $\int\omega_{t}\,\mathrm{d}x=0$ for example. This is also reflected by the parabolic maximum principle:
$\inf\omega_{t}\geq\inf\omega_{0}=\inf g_{0}\geq 0,$
which shows that the vorticity cannot change sign. We view it as an essential limitation for physical considerations, where one may generally face the $L^{1}$ vorticity function, while in this article we resolve this difficulty by allowing the circulations to be positive at some points and negative elsewhere.
1.1. Liouville Equation, BBGKY Hierarchy and Limit Equation
In this subsection, we present the corresponding Liouville (master) equation (forward Kolmogorov equation) and the BBGKY hierarchy of the SDE system (1.1) and the nonlinear Fokker-Planck equation of the limit system (1.3), since our method will always work on the PDE level. Hereinafter we keep the following notations:
$\displaystyle{\mathcal{M}}^{N}=(m_{1},\cdots,m_{N})\in{\mathbb{R}}^{N},\quad X^{N}=(x_{1},\cdots,x_{N})\in{\mathbb{R}}^{2N},$
$\displaystyle Z^{N}=({\mathcal{M}}^{N},X^{N})\in\mathbb{D}^{N}=({\mathbb{R}}\times{\mathbb{R}}^{2})^{N},\quad z=(m,x)\in\mathbb{D}={\mathbb{R}}\times{\mathbb{R}}^{2}.$
Denote by $G_{t}^{N}({\mathcal{M}}^{N},X^{N})$ the joint distribution of ${\mathcal{M}}^{N}$ and $X^{N}$ for the particle system at time $t$, with the factorized initial data:
$G_{0}^{N}(Z^{N})=g_{0}^{\otimes N}(Z^{N})=\prod_{i=1}^{N}g_{0}(z_{i}),$
which corresponds to the choice of i.i.d. initial data of the point vortex system (1.1): $\{\mathcal{M}_{i},X_{i}(0)\}_{1\leq i\leq N}$ are $N$ independent copies of $(\mathcal{M},X_{0})$ with law $g_{0}$. We also define the $k$-marginal of the joint law of the particle system by
$G_{t}^{N,k}(z_{1},\cdots,z_{k})=\int_{\mathbb{D}^{N-k}}G_{t}^{N}(Z^{N})\,\mathrm{d}z_{k+1}\cdots\,\mathrm{d}z_{N}.$
@@ 1.1. Liouville Equation, BBGKY Hierarchy and Limit Equation
We can relax the factorization assumption on the initial data $G_{0}^{N}$. For our purposes below, it is sufficient to assume that the marginals of all orders of the initial distribution $G_{0}^{N,k}$ are close to their factorized counterparts $g_{0}^{\otimes k}$ in the sense of relative entropy as below:
$H(G_{0}^{N,k}|g_{0}^{\otimes k})=\int_{\mathbb{D}^{k}}G_{0}^{k}\log\frac{G_{0}^{k}}{g_{0}^{\otimes k}}\,\mathrm{d}Z^{k}\lesssim\frac{k^{2}}{N^{2}}.$
We choose the simplest factorization assumption here, which corresponds to that the relative entropy $H(G_{0}^{N,k}|g_{0}^{\otimes k})$ vanishes, but our results can be easily applied to the above more general case.
Applying Itô formula to the particle system (1.1), one concludes that the joint law $G_{t}^{N}$ solves the following Liouville equation:
@@
We emphasize that $G_{t}^{N}$ shall be understood as the entropy solution of the Liouville equation (1.5). The definition of entropy solution is given as in Jabin-Wang [30]. For the sake of completeness, we state it as follows.
@@
A probability density function $G_{t}^{N}\in L^{1}(\mathbb{D}^{N})$ with $G_{t}^{N}\geq 0$ and $\int_{\mathbb{D}^{N}}G_{t}^{N}\,\mathrm{d}Z^{N}=1$ is called an entropy solution to the Liouville equation (1.5) in $[0,T]$, if $G_{t}^{N}$ solves (1.5) in the sense of distribution and for a.e. $t\in[0,T]$ it holds that
$\,\int_{\mathbb{D}^{N}}G_{t}^{N}\log G_{t}^{N}\,\mathrm{d}Z^{N}+\sigma\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{D}^{N}}\frac{|\nabla_{x_{i}}G_{s}^{N}|^{2}}{G_{s}^{N}}\,\mathrm{d}Z^{N}\,\mathrm{d}s$
$\leq$ $\,\int_{\mathbb{D}^{N}}G_{0}^{N}\log G_{0}^{N}\,\mathrm{d}Z^{N}.$
For the existence of such entropy solution, we also refer to [30].
@@
$+\frac{N-k}{N}\sum_{i=1}^{k}\int_{\mathbb{D}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G_{t}^{N,k+1}\,\mathrm{d}z_{k+1}=\sum_{i=1}^{k}\Delta_{x_{i}}G_{t}^{N,k}.$
Hereinafter, we assume that $1\leq k\leq N$ and use the convention that $G_{t}^{N,N+1}=0$. The above equation (1.6) describes the evolution of the $k$-marginal using the next order $(k+1)$-marginal, hence forms a hierarchy of PDEs, which is indeed the BBGKY hierarchy of the particle system.
As for the limit system (1.3), by applying Itô formula, we also obtain that the limit density $g_{t}(m,x)$ solves the following nonlinear Fokker-Planck equation
@@
After direct and simple computations, one can verify that the $k$-factorized law of $g_{t}(m,x)$, denoted by $g_{t}^{k}({\mathcal{M}}^{k},X^{k})$, solves the following PDE:
@@
Finally we define our normalized relative entropy between the joint distribution of the particle system and the $k$-factorized law of the limit system, which will be the major quantity that we shall estimate.
@@
The normalized relative entropy between two probability measures $F_{k}$ and $f_{k}$ on $\mathbb{D}^{k}$ is defined by
$H_{k}(F_{k}|f_{k})=\frac{1}{k}\int_{\mathbb{D}^{k}}F_{k}\log\frac{F_{k}}{f_{k}}\,\mathrm{d}Z^{k}.$
Hereinafter, we generally drop the index $t$ for simplicity. Also, since the variable ${\mathcal{M}}$ is independent of time $t$, if we let $f^{m}(x)=g(m,x)=\mathbb{P}(X|{\mathcal{M}}=m)$ be the conditional probability density of $X$ under ${\mathcal{M}}=m$ as in [17], we see that $f^{m}$ satisfies the same evolution equation (1.7) with $g$. Therefore, we shall not make the difference between $g(m,x)$ and $f^{m}(x)$ for calculations, and we assume that all derivatives of $g$ without index only act on the $x$ variable. In other words, we view $m$ as a fixed parameter.